OpenStudy (anonymous):

-x^2+3x+40 factoring quadratic expressions help please? i'v tried everything I know I cant figure it out I factored and the only multiples that get 3 is 5 and 8 but when I put them into the equation to check it I cant get it to come out right for nothing

OpenStudy (jango_in_dtown):

-x^2+3x+40 =-(x^2-3x-40)=-(x^2-2.x.3/2+9/4-9/4-40)=-{(x-3/2)^2-(13/2)^2}=(13/2)^2-(x-3/2)^2=(13/2+x-3/2)(13/2-x+3/2)=(5+x)(8-x)

OpenStudy (jango_in_dtown):

@_YouDontKnowHer_

OpenStudy (anonymous):

whoa that is a lot of stuff way different from how I got taught lol but thank you I will do my best to see what you did and figure it out

OpenStudy (jango_in_dtown):

Will I do in short??

OpenStudy (jango_in_dtown):

@_YouDontKnowHer_

OpenStudy (jango_in_dtown):

\[-x^2+3x+40=-x^2+8x-5x+40=-x(x-8)-5(x-8)=(-x-5)(x-8)=(x+5)(8-x)\]

OpenStudy (anonymous):

if you want to it's just the way we learned it this is how I do it -x^2+3x+40 factors of 40 are 1*40, 2*20, 4*10, and 5*8 which of the factors equal the middle number 3 8-5 is 3 so 8 and 5 are the only factors that will fit so always put the bigger number in the parenthesis first ex. -(x+8)-(x+5) drop down your first sign + then multiply positive times the other sign positive equals positive so your other sign is also positive that's how I got both my signs then this is where I got stuck checking it -(x+8)-(x+5) -x*-x= -x^2 -x*+5= -5 8*-x= -8x 8*5=40 so then that gives you the equation -x^2-5-8x+40 but -5-8x isnt a positive 3 so I dont know what I did wrong

OpenStudy (anonymous):

@jango_IN_DTOWN

OpenStudy (jango_in_dtown):

YOu saw my solution? Which line do you have a problem?

OpenStudy (jango_in_dtown):

you made a mistake here

OpenStudy (jango_in_dtown):

it will be -x(x-8)-5(x-8)

OpenStudy (jango_in_dtown):

@_YouDontKnowHer_